HUBBLE SPACE TELESCOPE DAILY REPORT #829 PERIOD COVERED: 0800 EST Daily Status Report as of 054/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC/PC 4207 (Photometric Calibration Monitor) Both the Planetary Cmaera and the Wide Field Camera was used to make ultraviolet flux observations of standard star DB+75D325 to obtain single chip readouts of W2 and P6 using wide band filters. The observations were executed as scheduled, and no problems were reported. 1.2 Completed PC 4714 (WF/PC Calibration: Internal Delta Flats) The Planetary Camera was used to obtain four internal flat observations. All four observations were received and displayed in the OSS. No problems were reported. 1.3 Completed Two Set of FGS 4064 (Crit. Test of Galactic Escape Velocity at R(Sun)) The Fine Guidance Sensors were used to make two sets of astrometric observations of the fast moving star G166-37. There were two recenterings during each series of astrometric observations. The last recentering occurred during an AOS and lasted for 17 seconds. 1.4 Completed Three Sets of WFC 4713 (WF/PC Calibration: Internal Delta Flats) The Wide Field Camera was used to make three internal flat observations. These images will be used for the restoration of Earth flats, which are altered due to the changes in quantum efficiency which occurs after the WFPC decontamination operations. All three images were received and displayed. They were reported to look normal. 1.5 Completed WFC 4105 (HST Medium Deep Survey: High Latitude) The Wide Field Camera was used to make two high latitude observations taken in parallel with the prime science observations. Both observations were received and displayed. they were reported to look normal. 1.6 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.7 Completed GHRS/2 4068 (GHRS Detector Calibration) The High Resolution Spectrograph was used to make one dark and two internal flat observations to check on the stability of the HRS detectors. All three observations were received and displayed. They all were reported to look okay. 1.8 Completed PC 3769 (Structure of the Inner Coma of Comet Chiron) A total of nine images were taken using the Planetary Camera and filter F555W to study the comet/object 2060 Chiron. This is to find the outer edge (i.e, the exopause) of the gravitationally bound dust cloud around the nucleus (i.e., the inner coma) and to look for indications of active regions within the inner coma. All nine of the images were received and displayed. They were reported to look normal. 1.9 Initiated FOS/BL 3545 (Evolution of the UV Spectra in Early Type Galaxies) The Faint Object Spectrograph was used to initiate ultraviolet observations of the early type galaxy 55W041. The observations are currently under way, and no problems have been reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 6 Successful : 6 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Significant HSTARs: None 2.4 Operations Note: A special FGS #2 analysis and review meeting to study the anomaly is currently under way at GSFC (Tuesday, Feb. 23, 1993). 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of FOS/BL 3545 (Evolution of Ultraviolet Spectra in Early Type Galaxies) 3.2 Complete WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) 3.3 Complete Two Sets of PC 4714 (WF/PC Calibration: PC Internal Delta Flats) 3.4 Complete WFC 4713 (WF/PC Calibration: Internal Delta Flats) 3.5 Complete PC 3769 (Spectra of Inner Coma of Comet Chrion) 3.6 Complete GHRS/2 3862 (Excitation of Atmospheres of Planetary Satellites) 3.7 Complete Two Sets of FOS/RD 4784 (Transition Comets: UV Search for OH) 3.8 Complete PC 3230 (PC PSF and Focus Monitoring 2) 3.9 Initiate HSP/UV1 4209 (HSP Detector Sensitivity Test) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-